# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ 2.3 Circuit Discovery
A circuit is a subgraph in a computational graph that is responsible for a certain behavior. Recent studies generally use causal intervention for circuit discovery. (Meng et al., 2023) proposed activation patching to identify activations in a model relevant to the output, while (Nanda, 2023) proposed attribution patching to accelerate it. (Wang et al., 2022; Conmy et al., 2023) focused on edges and proposed path patching. Others optimized this technique from various aspects (Syed et al., 2023; Kramár et al., 2024).
2.4 Intrinsic Dimension
(Li et al., 2018) first proposed the concept of intrinsic dimension to describe the minimum number of dimensions needed for solving a specific problem. (Aghajanyan et al., 2020) provided an analysis into fine-tuning with this concept and proved the feasibility of effective fine-tuning. Inspired by previous works, (Hu et al., 2021) proposed LoRA as a parameter-efficient fine-tuning method, which is of great significance for fine-tuning large language models.
3 The Theory of Node Redundancy
In this section, we introduce the concept of node redundancy as an extension of the intrinsic dimension, together with a method for detecting it, which can also be used to explain the idea behind circuit discovery from a new perspective.
Since the intrinsic dimension describes the redundancy in parameters, our goal is also to find out the redundant parameters in a model. If we inspect a representation $H=({h_{1},h_{2},...,h_{m}})^{\top}\in\mathbb{V}^{D}(D=m)$, then we can denote the parameter matrix to examine as $W^{pre}\in\mathbb{R}^{m\times n}$, which maps an input from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$.
3.1 Representation Redundancy
Assumption 3.1.
(Feature Redundancy) Suppose all features represented in a neural network is $\mathcal{F}$. Given a task $T$ with $\mathcal{X}=\{x_{1},x_{2},...,x_{t}\}\sim\mathcal{D}_{T}$ which consists of $t$ samples that follows a specific data distribution $\mathcal{D}_{T}$, the features for representing $\mathcal{X}$ is $\mathcal{F}_{T}\subsetneqq\mathcal{F}$.
Remark 3.2.
3.1 means that when conditioned on a task $T$, the number of features needed is smaller than that of the elements in the universal set $\mathcal{F}$. The assumption follows the concept of feature sparsity from (Elhage et al., 2022) that many features do not frequently appear.
Definition 3.3.
(Semantic Preservation) Consider a feature $f$ with its feature vector $\mathbf{v_{f}}=\sum_{i=1}^{m}c_{i}\mathbf{e_{i}}$. If we set the coefficients in $\mathbf{v_{f}}$ which correspond to a set of bases $\mathbf{E_{r}}$ to zero, then we can get a new vector $\tilde{\mathbf{v_{f}}}$. If the cosine similarity $cos(\mathbf{v_{f}},\tilde{\mathbf{v_{f}}})=\frac{\mathbf{v_{f}}\cdot\tilde{\mathbf{v_{f}}}}{||\mathbf{v_{f}}||\cdot||\tilde{\mathbf{v_{f}}}||}>\eta$ $(\eta\in(0,1])$, then we say $\tilde{\mathbf{v_{f}}}$ has preserved the semantic information in $f$.
Remark 3.4.
It is a nice property akin to the robustness of population encoding to noise. Since information is encoded across many cells, a perturbation on a few cells will not destroy the representation (Pouget et al., 2000).
Definition 3.5.
(Dimension Redundancy) Given a task $T$ with a number of features $\mathcal{F}_{T}$. Let’s consider a single basis $\mathbf{e_{i}}$ in the representation space $\mathbb{V}^{m}$. For any feature $f$ in $\mathcal{F}_{T}$, we set the coefficient $c_{i}$ in $\mathbf{v_{f}}$ to zero and get the deformed feature vector $\tilde{\mathbf{v_{f}}}$. If $\mathbf{e_{i}}$ satisfies:
$min[cos(\mathbf{v_{f_{1}}},\tilde{\mathbf{v_{f_{1}}}}),...,cos(\mathbf{v_{f_{|\mathcal{F}_{T}|}}},\tilde{\mathbf{v_{f_{|\mathcal{F}_{T}|}}}})]>\eta$
then $\mathbf{e_{i}}$ is a redundant dimension for representing features in task $T$. Note that $\eta\in(0,1]$ serves as the lower bound for judging whether $\tilde{\mathbf{v_{f}}}$ preserves the information in $f$.
Proposition 3.6.
Consider the setup of Definition 3.5. If $\mathbf{e_{i}}$ satisfies:
$max(|\mathbf{v_{f_{1}}}\mathbf{e_{i}}|,|\mathbf{v_{f_{2}}}\mathbf{e_{i}}|,...,|\mathbf{v_{f_{|\mathcal{F}_{T}|}}}\mathbf{e_{i}}|)<\xi$
where $\xi=\sqrt{1-\eta^{2}}$, then we say $\mathbf{e_{i}}$ is redundant for representing features in task $T$.
Remark 3.7.
This is obvious from Definition 3.5. Note that $|\mathbf{v_{f_{j}}}\mathbf{e_{i}}|$ is the projection of $\mathbf{v_{f_{j}}}$ on direction $\mathbf{e_{i}}$. Higher $|\mathbf{v_{f_{j}}}\mathbf{e_{i}}|$ means higher dependency of $\mathbf{v_{f_{j}}}$ on $\mathbf{e_{i}}$. The redundancy probability $P_{r}$ of a single basis $\mathbf{e_{i}}$ can be written as:
$P_{r}=\prod_{j=1}^{|\mathcal{F}_{T}|}P(|\mathbf{v_{f_{j}}}\mathbf{e_{i}}|<\xi)$
The equation tells us that the probability a basis $\mathbf{e_{i}}$ is redundant will decrease with the increasing number of features. On the contrary, the fewer features related to a specific task, the higher the likelihood that a certain basis is redundant, leading to more redundant dimensions in the representation.
Definition 3.8.
(Representation Redundancy) Given a specific task $T$ with features $\mathcal{F}_{T}$, we choose a set of bases $\mathbf{E_{r}}\subset\mathbf{E}$. For each feature $f\in\mathcal{F}_{T}$, we set the coefficients in its feature vector $\mathbf{v_{f}}$ which correspond to the bases in $\mathbf{E_{r}}$ to zero and get the deformed feature vector $\tilde{\mathbf{v_{f}}}$. Given a threshold $\tau_{r}$ as the tolerance of deformation. If $\mathbf{E_{r}}$ satisfies:
$\sum_{j=1}^{|\mathcal{F}_{T}|}max[(\eta-cos(\mathbf{v_{f}},\tilde{\mathbf{v_{f}}})),0]<\tau_{r}$
then $\mathbf{E_{r}}$ is a set of redundant bases for representing features $\mathcal{F}_{T}$, and the representation redundancy in $\mathbb{R}^{m}$ is defined under the choice of $\mathbf{E_{r}}$. If $|\mathbf{E_{r}}|=r$, then the value of representation redundancy is equal to $r$.
Remark 3.9.
The definition of representation redundancy is actually an extension of Definition 3.5 from one-dimension to multi-dimension. The difference $\eta-cos(\mathbf{v_{f}},\tilde{\mathbf{v_{f}}})$ is used to measure the degree of deformation of feature vectors. If the total deformation in feature vectors is accepted under the threshold $\tau_{r}$, then the influence from the removal of $\mathbf{E_{r}}$ on the semantic representation of $\mathcal{F}_{T}$ can be ignored, and the number of useful dimensions for representing $\mathcal{F}_{T}$ is $m-r$.
3.2 Forward Redundancy
If the representation $H=(h_{1},h_{2},...,h_{m})^{\top}$ is followed by another weight matrix $W^{post}\in\mathbb{R}^{l\times m}$, then the following representation is $H^{\prime}=W^{post}H\in\mathbb{V}^{D^{\prime}}(D^{\prime}=l)$. To investigate into the influence of $H$ on $H^{\prime}$, we split $H^{\prime}$ into the sum of influences from the dimensions in $H$:
$W^{post}H=\sum_{i=1}^{m}W_{:,i}^{post}h_{i}$ (2)
From (2), we find that the influence from $h_{i}$ to $H^{\prime}$ is subject to $W_{:,i}^{post}$. If the value of $W_{:,i}^{post}$ is small, then the influence from $h_{i}$ will be weakened, leading to a reduced impact on the final result in the follow-up calculations.
Definition 3.10.
(Forward Redundancy) Given a representation $H$ and the follow-up parameters $W^{post}$, we inspect the $i$-th dimension in $H$. Given a lower bound $\tau_{f}<0$ for the influence of $H$ in the forward propagation.If
$||W_{:,i}^{post}h_{i}||<\tau_{f}$
then we say the dimension $i$ in $\mathbb{R}^{m}$ is redundant in the forward propagation. If the number of redundant dimensions under this condition is $r$, then the value of forward redundancy is equal to $r$.
3.3 Node Redundancy
3.3.1 The definition of node redundancy
The discussions in previous sections focus on the real computations inside a model. In this section, we regard the whole model as a DAG composed of nodes and edges. The node could be a single neuron or an attention head in Transformer, based on the level of granularity.
To determine whether a node is redundant or not, a method is needed for measuring the importance of that node. Inspired by the idea of causal intervention in the area of causal inference, we provide the following definition:
Definition 3.11.
(Node Redundancy) Given a model $M$ and its computational graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in which $\mathcal{V}$ and $\mathcal{E}$ represent the nodes and edges in $\mathcal{G}$ respectively, consider a node $n\in\mathcal{V}$. For task $T$, a batch of inputs $\mathcal{X}=\{x_{i}\}_{i=1}^{p}\sim\mathcal{D}_{T}$ is fed into the model and the activation of $n$ is saved.
For each input $x_{i}$, if we replace $n(x_{i})$ which is the value of node $n$ with another value $n(x_{i}^{\prime})$ corresponding to the corrupted input $x_{i}^{\prime}$ while keeping other activations in the forward propagation unchanged, we can get another output $M(x_{i}^{\prime})$. Given a metric $f$ for measuring the difference between two output distributions caused by $n$ and a threshold $\tau$, we can calculate the contribution $c_{i}(n)$ of the node $n$ as:
$c_{i}(n)=f(n;M(x_{i}),M(x_{i}^{\prime}))$
If the contribution of $n$ satisfies:
$\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|c_{i}(n)|\big{]}<\tau$
then $n$ is a redundant node in $\mathcal{G}$. If the redundant nodes in terms of task $T$ in model $M$ is $\mathcal{N}$, then the value of the intrinsic dimension at the node level is $|\mathcal{V}-\mathcal{N}|$.
3.3.2 The detection of node redundancy
With the idea of intervention from causal inference, we adopt the concept of indirect effect (Pearl, 2001) to estimate the contribution of a node $n$ to the output as:
$IE(n;x)=\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n\leftarrow n(x^{\prime})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}$
in which $IE$ is the indirect effect from node $n$ to the final output of the model, and $\mathcal{L}_{m}$ is a metric for measuring the final output. We use the do-calculus notation (Neuberg, 2003) to express the intervention behavior, which is also called patching. The metric $IE$ measures the effect of an intermediate variable in a causal graph to the final behavior of the model, which is widely used in causal inference. If $\mathcal{L}_{m}$ is viewed as a function of $n$, we can apply a first-order Taylor expansion to $IE$ at $n=n(x^{\prime})$ for approximation. Thus $IE(n;x)$ can be simplified as
$\begin{split}&\mathcal{L}_{m}[M(x)]\!+\![n(x^{\prime})\!-\!n(x)]^{\top}\nabla_{n}\mathcal{L}_{m}[M(x)]\!-\!\mathcal{L}_{m}[M(x)]\\ =&[n(x^{\prime})\!-\!n(x)]^{\top}\nabla_{n}\mathcal{L}_{m}[M(x)]\big{|}_{n(x)}\end{split}$
If zero ablation is used which means $n(x^{\prime})$ is set to zero, then the contribution of node $n$ can be written as:
$c(n)\approx\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}n\nabla_{n}\mathcal{L}_{m}[M(x)]\big{]}$ (3)
The IE in equation (3) consists of two parts: (i) $n$ the value of node $n$ on the clean input $x$ and (ii) $\nabla_{n}\mathcal{L}_{m}[M(x)]$ the derivatives of the patching metric $\mathcal{L}_{m}$ with respect to node $n$. The two parts correspond to the representation redundancy and the forward redundancy respectively. To gain an intuitive understanding of this, let’s consider a representation in $\mathbb{V}^{m}$ which consists of a set of nodes $\mathcal{N}=\{n_{1},n_{2},...,n_{m}\}\subset\mathcal{V}$ at the neuron level and then focus on the node $n_{i}$:
  • The first part of (3) directly serves as an indicator for representation redundancy since the magnitude of a neuron is a good indicator for dimension redundancy. According to (1), the smaller the value of a neuron is, the higher the likelihood that the neuron corresponds to a redundant dimension, since the activation of a feature $f$ on a basis $\mathbf{e_{i}}$ is directly proportional to the projection of its feature vector $\mathbf{v_{f}}$ on $\mathbf{e_{i}}$.
    One thing to note is the case where two features $f_{1}$ and $f_{2}$ are almost in opposite directions and the activations in the shared direction may cancel out each other, which results in a situation that the value of a neuron seems to be small though both of the two features fire on it. While this case may be tricky, (Elhage et al., 2022) show that models prefer to represent anti-correlated features in opposite directions, which means $f_{1}$ and $f_{2}$ may never co-occur at the same time, so the probability of the above case occurring is very small.
  • As for the second part, according to the chain rule during the back propagation stage, it can be written as:
    $\nabla_{n_{i}}\mathcal{L}_{m}=\sum_{j=1}^{l}\nabla_{n_{j}}\mathcal{L}_{m}\cdot\nabla_{n_{i}}n_{j}$
    where the node $n_{j}$ in the follow-up vector space $\mathbb{V}^{l}$ is
    $n_{j}=\sum_{i=1}^{m}w_{ji}^{post}n_{i}$
    Thus
    $\nabla_{n_{i}}\mathcal{L}_{m}=\sum_{j=1}^{l}\nabla_{n_{j}}\mathcal{L}_{m}\cdot w_{ji}^{post}$ (4)
    From equation (4), it can be seen that the second part in (3) is closely related to the follow-up calculations as discussed before in Section 3.2, hence serving as an indicator for forward redundancy.
As discussed above, we can see that the metric for judging the redundancy of a node is a combination of the representation redundancy and the forward redundancy. Therefore the essence of the intrinsic dimension at the node level can be decomposed and understood from the above two aspects.
Note that Section 3.3.2 is also adopted in circuit discovery, e.g., attribution patching (Nanda, 2023). Thus the process of circuit discovery can be viewed as a process of identifying the two types of redundancy and finding out the intrinsic dimensions in terms of a task. Each of the intrinsic dimensions corresponds to a node in the computational graph, and the number of the nodes in the circuit is expected to be equal to the dimensionality of the solution space. Given a specific task, the co-occurrence of the representation redundancy and the forward redundancy is the necessary and sufficient condition of the node redundancy in the graph.

### v2

@@ 2.2 Mechanistic interpretability for learning
Mechanistic interpretability aims to understand the computational mechanisms of a model [53]. Existing works can be divided into static research on the trained model and dynamic research during training according to whether the model is in an inference state or a training state. In the static research, circuit discovery is a widely used technique that aims to look for the subgraph in a computational graph that acts as an algorithm implemented in a model for a certain behavior. Meng et al. [35] proposed activation patching to identify activations relevant to the output, while Nanda [37] proposed attribution patching to accelerate it. Wang et al. [62], Conmy et al. [9] focused on edges and proposed path patching. Others optimized this technique from various aspects [56, 27, 34, 2, 32]. Despite their success, current studies are limited to discover the circuit of an existing model behavior of a trained model in a static style, while our method is able to form a new circuit of a non-existent capability through iterations in a dynamic mode during training.
Compared to static research, there is fewer research on learning dynamics. For phase transition, Nanda et al. [38] delved into the study of grokking, while Olsson et al. [42] analyzed the emergency of induction heads for in-context learning. For fine-tuning mechanisms, Prakash et al. [46] focused on entity linking, Jain et al. [23] focused on compiled models and probabilistic context-free grammars, Wang et al. [63] focused on math, while Lee et al. [28] focused on DPO for toxicity reduction. Kotha et al. [26] explored the catastrophic forgetting. Wu et al. [64] proposed ReFT that learns an intervention on the representations for efficient fine-tuning. Parallel to our work, Ren and Sutherland [50] proposed a framework for learning dynamics by decomposing the change of a model’s prediction. However, current studied are limited to a single scenario with various analytic methods, while there lacks a unified method for the study of learning dynamics on common tasks, and no attempt was made to actively introduce interpretability into fine-tuning for mechanistic study.
Unlike prior works, we successfully transfer the concept of intrinsic dimensionality from the parameter space to the representation space, and dynamically integrate circuit discovery into fine-tuning as a heuristic method for the study of learning dynamics for the first time, providing new insights for the study of learning dynamics from the mechanistic view.
3 The theory of node-level intrinsic dimensionality
3.1 Background
As discussed in Section 2.1, intrinsic dimensionality implies the redundancy in parameters. Measuring the exact intrinsic dimensionality is computational intractable. Therefore heuristic methods are adopted to measure the upperbound of it. Considering a set of parameters $\theta^{D}=[\theta_{0},...,\theta_{m}]$ and an objective function $f(\cdot,\theta)$, Li et al. [31] leveraged a re-parameterization method to optimize only the parameters $\theta^{d}$ in a subspace $\mathbb{R}^{d}(d<D)$ via a linear transformation with a projection matrix $P$:
$\theta^{D}=\theta_{0}^{D}+P(\theta^{d})$ (1)
If a satisfactory solution is reached, then the dimensionality of that subspace is the intrinsic dimensionality. In practice, the heuristic method requires searching over various $d$, optimizing the parameters $\theta^{d}$ and selecting the smallest $d$ that could reach a satisfactory result.
3.2 Redefine intrinsic dimensionality
In this section, we view the model $M$ as a computational graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}$ and $\mathcal{E}$ represent the nodes and edges in $\mathcal{G}$ respectively. Each node is a vector $n=(h_{1},...,h_{N})^{\top}\in\mathbb{R}^{N}(1\leq N\leq D)$ that could be the activation of a neuron, a group of neurons or a representation in a $D$-dimensional representation space $\mathbb{V}^{D}$, based on granularities of interest. The edges describe the information flows between the nodes, i.e., where the output of an upstream node goes and where the input of a downstream node is from.
Note that it is possible to set a node to a full layer, or a latent which is a sparse feature in dictionary learning, while here we require the size of a node to be no more than the dimensionality of $\mathbb{V}^{D}$ to cover most of the cases in practice and for convenience of discussion. Also, an edge is not necessary to follow the real structure of a model. It could be a virtual connection between non-adjacent nodes.
Since the decomposition of a model in mechanistic interpretability often looks at activations or latents [53], we look for the intrinsic dimensions indirectly via the activations in the representation space. Inspired by prior works on circuit discovery, we believe that the learning process of a model is to dynamically identify the redundancy in a computational graph and optimize the subgraph of a task.
Definition 1: intrinsic dimensionality
Given the computational graph $\mathcal{G}$ of a model and a task $T$, the intrinsic dimensionality is the number of nodes in the minimal subgraph $\mathcal{C}^{*}=\{\mathcal{V}_{T}^{*},\mathcal{E}_{T}^{*}\}\subset\mathcal{G}$ on which the optimization of the objective function for task $T$ could reach a predefined precision level.
The newly defined intrinsic dimensionality, namely the node-level dimensionality, describes the redundancy in parameters indirectly with the "intrinsic nodes" in the minimal subgraph. Under this new definition, the parameters necessary for a task $T$ refer to those where the intrinsic nodes $\mathcal{V}_{T}^{*}$ are derived from. For example, if a parameter matrix $W\in\mathbb{R}^{m\times n}$ maps an input $x\in\mathbb{R}^{n}$ to a representation $H\in\mathbb{R}^{m}$ that is an intrinsic node, then $W$ is what matters to task $T$. A theoretical explanation of this is shown in Section 3.4 and detailed in Appendix E.
3.3 Estimating intrinsic dimensionality
Different from prior works that measure the intrinsic dimensionality in a subspace, we choose to perform the estimation in a subgraph. Similar to Li et al. [31], we also adopt a heuristic method to estimate the intrinsic dimensionality through training. The difference is that instead of optimizing via re-parameterization, we directly search for the intrinsic nodes in the computational graph through iterations. Inspired by the idea of causal intervention for the structural causal model [39], we adopt the concept of indirect effect [44]. Specifically, let’s consider a node $n\in\mathcal{V}$. For the task $T$, a batch of inputs $\mathcal{X}=\{x_{i}\}_{i=1}^{p}\sim\mathcal{D}_{T}$ is fed into the model $M$. For each input $x$, if we replace $n(x)$ that is the activation of the node $n$ with another value $n(x^{\prime})$ from a corrupted input $x^{\prime}$ while keeping other activations unchanged, we can get another output $M(x^{\prime})$. Given a metric $\mathcal{L}_{m}$ for measuring the output, the contribution $c(n)$ of the node $n$ in terms of the input $x$ is its indirect effect on the output:
$c(n)=\Big{|}\mathcal{L}_{m}\big{[}M\big{(}x\,|\,do\big{(}n\leftarrow n(x^{\prime})\big{)}\big{)}\big{]}-\mathcal{L}_{m}\big{[}M(x)\big{]}\Big{|}$ (2)
The do-calculus notation [39] is used to express the intervention behavior in Equation (2). Given a lowerbound $\tau$, If the contribution of $n$ satisfies: $\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}c_{i}(n)\big{]}>\tau$, then $n$ is an intrinsic node. During optimization, we dynamically prune the graph to encourage exploration over the computational graph. When a satisfactory precision of optimization is reached, then the intrinsic dimensionality is the minimum number of intrinsic nodes $|\mathcal{V}_{T}^{*}|$ in the final subgraph for the task $T$.
3.4 Further discussions on the node-level intrinsic dimensionality
What does the indirect effect imply?
As discussed in Section 3.3, the indirect effect is adopted to measure the importance of a node in a causal manner. Since intrinsic dimensionality originates from the parameter redundancy in a model [29, 1], while a node is defined over the activations, how can we describe parameter redundancy through the lens of activations, and what exactly are the redundancies in a representation? To gain a deeper understanding of it, we provide a theoretical analysis detailed in Appendix E. In short, we decompose Equation (2) into two parts by applying a first-order Taylor expansion to the right side of it at $n=n(x)$ for approximation:
(2$)\!\approx\!\Big{|}\mathcal{L}_{m}[M(x)]\!+\![n(x^{\prime})\!-\!n(x)]^{\top}\nabla_{n}\mathcal{L}_{m}[M(x)]\!-\!\mathcal{L}_{m}[M(x)]\Big{|}\!=\!\Big{|}[n(x)\!-\!n(x^{\prime})]^{\top}\!\nabla_{n}\mathcal{L}_{m}[M(x)]\Big{|}$ (3)
In practice, the intervention on the node $n$ is performed using the averaged result of all other nodes. Therefore $n(x)-n(x^{\prime})$ can be viewed as a centered activation $n(x)^{c}$. Thus Equation (3) consists of two parts: (i) the first part $n(x)^{c}$ is an indicator for representation redundancy, which describes the redundancy of a node for representing an input; (ii) the second part $\nabla_{n}\mathcal{L}_{m}[M(x)]$ is an indicator for forward redundancy, which describes the redundancy of a node in the forward propagation. Thus the estimation of the intrinsic dimensionality can be converted to the process of redundancy detection in a graph. The definitions of the redundancies as well as the proof are detailed in Appendix E.
The minimum description length (MDL) in a computational graph.
The MDL principle [51] encourages us to minimize the combined cost of encoding the model and the data given the model. While the parameters for the task $T$ may not be decoupled from others and are perhaps distributed over the whole parameter space $\mathbb{R}^{D}$, prior works on intrinsic dimensionality used a reparameterization trick to separate the intrinsic dimensions $\theta^{d}\in\mathbb{R}^{d}$ out from the original parameter space. This kind of measurement aims to look for the upperbound of the MDL in terms of a task.
While we aim to (i) find out the minimal subgraph $\mathcal{C}^{*}$ from a computational graph $\mathcal{G}$ and (ii) interpret the learning process of a model based on the subgraph, there is a need to balance the minimum of the required parameters and the convenience for interpretable graph structure. Since we want to keep the original topological structure of the computational graph for further analyses, we directly delve into the original graph to inspect the redundancy, instead of adding extra structures like linear projections. Therefore, the minimal subgraph does not mean the minimum parameters in a strict sense.
However, the functionalities of the nodes in a computational graph are thought to be decoupled from each other in the mechanistic study. Braun et al. [3] argued that one of the properties of a good decomposition of a model is the minimality that the decomposition should use as few components as possible to represent a model behavior of interest, which also follows the MDL principle. Therefore, our definition holds for the MDL principle in the sense of the minimality in mechanistic interpretability. Different from prior works, we do not require a strict estimation of the intrinsic dimensionality, but trying to take into account both the minimality of the graph and the feasibility for interpretability.